The balance sheet for Stud Clothiers is shown next. Sales for the year were $3,990,000, with 75 percent of sales sold on credit. STUD CLOTHIERS Balance Sheet 20X1 Assets Liabilities and Equity Cash Accounts receivable Inventory Plant and equipment $74,000 Accounts payable % 305,000 153,000 147,000 100,000 150,000 229,000 1,084,000 287,000 310,000 413,000 Accrued taxes Bonds payable (long-term) Common stock Paid-in capital Retained earnings Total assets $ 1,084,000 Total liabilities and equity Compute the following ratios: (Use a 360-day year. Do not round intermediate calculations. Round your answers to 2 decimal places. Input your debt-to-total assets answer as a percent rounded to 2 decimal places.) a. Current ratio b. Quick ratio c. Debt-to-total-assets ratio d Asset turnover e. Average collection period times times times days
Answer #1

Solution:

The total assets and liability are given and we need to identify the current assets and current liability to calculate the ratios.

Current asset = Cash + Account receivable + inventory = 74,000+287,000+310.000 = 671000

Current liability = Account payable+ Accrual tax = 305,000+153,000 = 458,000

A) Current ratio : This is a liquidity ratio that measures the short term liquidity of the company.

Current ratio = Current asset / Current liability = 671,000/458,000 = 1.47

B) Quick ratio = (Current asset - inventory) / current liability = 671,000-310,000)/458,000 = 0.79

C) Debt to total asset ratio = Debt / Total asset

Debt =147,000, total asset = 1,084,000

Debt to total asset ratio = Debt / Total asset = 147,000/1,084,000 = 13.56%

D) Asset turnover = Sales / Total asset

Sales = $3,999,000

Asset turnover = Sales / Total asset = $3,999,000 / 1,084,000 = 3.69

E) Average collection period = 360 / Account payable turnover

Account payable turnover = Credit sales / Account payable

Credit sales = 75% * 3,999,000 = 2,999,250

Account payable = 305,000

Account payable turnover = 2,999,250 / 305,000 =9.834

Average collection period = 360/9.834 =36.61 Days